Stewart created a beer process involving three departments the costs related to department B were as follows:

Category Units    % Material % Conversion costs
Beginning Inv 25000    20    2100 80 3000
Started 100000 285000    430000

Ending Inventory 50000    30 40
what is the cost per equivalent unit for material costs? 3.16 or 3.0 or None of the Above.

Pepper Company uses the weighted-average method in its process costing system. The following information was available for one of its processing departments:
• The beginning works in process inventory consisted of 16,000 units, which were 75% complete with respect to conversion costs.
• The department started 59,000 units in production during the current month.
• A total of 60,000 units were completed and transferred out of that department during the current month.
• The ending work in process inventory was 55% complete with respect to conversion costs and 35% for material

What is the number of equivalent units for conversion costs? 63200 or 65250 or 68250

At an activity level of 4,000 machine-hours in a month, Curt Corporation’s total variable production engineering cost is $154,200 and its total fixed production engineering cost is $129,000. If the activity level was increased to 4300 machine hours per month Assume that this level of activity is within the relevant range.

What is the variable cost per unit at 4300 hours? 38.5 or 6.3 or 7.0
Answer #1

1) Units completed and transferred= Beginning work in process+Units started into production-Ending work in process

= 25000+100000-50000= 75000

Calculation of Equivalent units for Materials

Materials
Units completed and transferred 75000
Ending work in process (50000*30%) 15000
Equivalent units for Materials 90000

Calculation of Cost per equivalent unit for material costs

Materials
Cost of beginning work in process $2100
Cost incurred during the period 285000
Total costs $287100
Equivalent units for Materials 90000
Cost per equivalent unit $3.19

So, the answer is None of the above

2) Ending work in process= Beginning work in process+Units started into production-Units completed and transferred out

= 16000+59000-60000= 15000

Conversion Costs
Units completed and transferred 60000
Ending work in process (15000*55%) 8250
Equivalent units for conversion costs 68250

So, the answer is 68250

3) Variable cost per unit= Total variable costs/Machine hours

= $154200/4000= $38.55

Variable cost per unit remains same irrespective of the number of machine hours

So, the answer is $38.5